# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
where the last equality is due to $[\bm{H}_{l}^{(i,1)},\dots,\bm{H}_{l}^{(i,p+1)}]=\bm{U}_{l-1}^{(i)^{T}}\bm{V}_{l}$ with $\bm{U}_{l-1},\ \bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$, and last inequality follows from $\|\bm{H}_{l}^{(i,j)}\|\leq 1$ for all $l\in[L]$ and $i,j\in[p+1]$. Therefore, we obtain
$\sum_{i=1}^{p+1}\|\bm{H}_{l}^{(i,i)}-\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}\|_{F}^{2}\leq\sum_{i=1}^{p+1}\sum_{j=1,j\neq i}^{p+1}\|\bm{H}_{l}^{(i,j)}\|_{F}^{2}.$
This, together with (129) and $\bm{T}_{l}^{(i)}=\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}$, yields
$\|\bm{H}_{l}-\mathrm{BlkDiag}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)})\|_{F}^{2}\leq\frac{81\sigma_{\max}^{*2}}{16\delta_{\sigma}^{2}\lambda^{2}\sigma^{*2}_{\min}}\|\nabla G(\bm{W})\|_{F}^{2},$
which directly implies (68). ∎
Now, we prove two key inequalities in Lemma 4, which will be frequently used in the subsequent analysis.
Proof of Lemma 4.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. It follows from Lemma 3 and 2 that (57)-(60) and (63) holds. Recall that (66) denotes an SVD of $\bm{W}_{l}$ for each $l\in[L]$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkDiag}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$ for each $l=2,\dots,L$. This, together with (68), implies
@@
where $c_{1}$ is defined in (64). Using the SVD (66), we have
$\left\|(\bm{W}_{L}\bm{W}^{T}_{L})^{L-1}\bm{W}_{L}-\sqrt{\lambda}\bm{Y}\bm{W}_{L-1:1}^{T}+\lambda\bm{W}_{L}\right\|_{F}$
$=$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}+\lambda\bm{\Sigma}_{L}\right\|_{F}$
$\geq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}(\bm{H}_{2}-\bm{T}_{2})\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}$
$-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}\bm{T}_{2}\bm{\Sigma}_{2}^{T}(\bm{H}_{3}-\bm{T}_{3})\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}-\dots-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\left(\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right)\bm{\Sigma}_{L-1}^{T}(\bm{H}_{L}-\bm{T}_{L})\right\|_{F}$
@@
where the first inequality uses the triangle inequality and the last inequality follows from (57) in Lemma 3, $\|\bm{H}_{l}\|=1$ and $\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (130). This, together with (131), implies
$\ \left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}$
@@
According to the block structure of $\bm{\Sigma}_{l}$ in (66) and $\bm{T}_{l}=\mathrm{BlkDiag}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$, we compute
$\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}=\mathrm{BlkDiag}\left(\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(1)}\bm{T}_{l+1}^{(1)},\dots,\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p)}\bm{T}_{l+1}^{(p)},\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p+1)^{T}}\bm{T}_{l+1}^{(p+1)}\right).$
@@
$\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\|_{F}^{2}$ $=\sum_{j=1}^{g_{i}}\left(\sigma_{j}(\bm{\Sigma}_{l}^{(i)})-\sigma_{j}(\bm{\Sigma}_{L}^{(i)})\right)^{2}=\sum_{j=1}^{g_{i}}\left(\sum_{k=l}^{L-1}\left(\sigma_{j}(\bm{\Sigma}_{k}^{(i)})-\sigma_{j}(\bm{\Sigma}_{k+1}^{(i)})\right)\right)^{2}$
$\leq g_{i}(L-1)^{2}\left(\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}\right)^{2},$
which implies
$\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\|_{F}\leq\frac{3\sqrt{2g_{\max}}L\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}.$ (133)
We claim that it holds for each $l\in[L]$ and $i\in[p]$ that
$\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\|_{F}\leq\frac{\eta_{1}}{\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},$ (134)
where $\eta_{1}$ is defined in (73). Using the triangular inequality, we obtain
$\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\ \sum_{j=2}^{L}\left\|\underbrace{\left(\prod_{l=1}^{L-j}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}-\left(\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\right)\bm{\Sigma}_{L}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{j-2}}_{\bm{R}_{j}}\right\|_{F}.$ (135)
@@
$\leq(j-1)\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{L-2}\frac{\eta_{1}}{\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F},$
where the first inequality uses (i) of Lemma 3 and $\|\bm{T}_{l+1}^{(i)}\|=1$ for all $l$ and $i$, the second one is due the triangular inequality, and the last one follows from (134). This, together with (4.2), implies
@@
$\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)(\bm{\Sigma}_{L}^{(i)})^{L-1}\right\|_{F}+\left\|(\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{2}^{(i)}\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+$
$\ \left\|\bm{\Sigma}_{L}^{(i)}\bm{T}_{2}^{(i)}(\bm{\Sigma}_{2}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{3}^{(i)}\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+\cdots+\left\|\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}(\bm{\Sigma}_{L-1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{L}^{(i)}\right\|_{F}$
@@
where the first inequality uses the triangular inequality and the last inequality follows from (136), (57) in Lemma 3, $\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (133). This, together with (72), yields
$\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkDiag}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
@@
$\ \left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\mathrm{BlkDiag}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
$\leq$ $\ \left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}+$
$\quad\sqrt{\lambda}y_{1}\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkDiag}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}\leq c_{2}\|\nabla G(\bm{W})\|_{F},$
where $c_{2}$ is defined in (74) and the last inequality follows from (4.2) and (4.2). Applying the same argument to $\nabla_{\bm{W}_{1}}G(\bm{W})$, we obtain (69).
The rest of the proof is devoted to proving the claim (134). According to (66) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l=2,\dots,L$, we have for each $i\in[p]$,
$\|\bm{W}_{l}^{T}\bm{W}_{l}-\bm{W}_{l-1}\bm{W}_{l-1}^{T}\|_{F}=\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{H}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\|_{F}$
$\geq$ $\|\bm{T}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{T}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\|_{F}-2\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\|_{F}$
$\geq$ $\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{l}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}-2\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\|_{F}$
$\geq$ $\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}-2\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\|_{F}-\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l}^{(i)})^{2}\|_{F}-\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\|_{F}$
$\geq$ $\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}-2\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\|_{F}-3\sigma^{*}_{\max}\left(\|\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{l}^{(i)}\|_{F}+\|\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{l-1}^{(i)}\|_{F}\right),$
where the second inequality follows from $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and the last inequality follows from (57). This together with (59), (130), and (133), yields that for each $l=2,\dots,L$ and $i\in[p]$, we have
$\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}\leq\eta_{1}\|\nabla G(\bm{W})\|_{F}.$ (138)
For each $i\in[p]$, using Weyl’s inequality yields for each $j\in[g_{i}]$,

### v3

@@ 3.2.1 Preliminary Results
where the last equality is due to $\bm{H}_{l}^{(i,1)},\dots,\bm{H}_{l}^{(i,p+1)}]=\bm{U}_{l-1}^{(i)^{T}}\bm{V}_{l}$ with $\bm{U}_{l-1},\ \bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$, and last inequality follows from $\|\bm{H}_{l}^{(i,j)}\|\leq 1$ for all $l\in[L]$ and $i,j\in[p+1]$. Therefore, we obtain
[…]
This, together with the definition of $\bm{H}_{l}$, (53) and $\bm{T}_{l}^{(i)}=\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}$, yields (48). ∎
Next, using the above proposition, we show that for weight matrices in the neighborhood of the set of critical points, the singular matrices $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{L}$ satisfy the following spectral inequalities.
Lemma 3.
Consider the setting of Lemma 2. Suppose that there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48) holds for all $l=2,\dots,L$. Let $c_{1}$ be defined in (42) and
$\textstyle\bm{A}_{i}:=\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{1}^{(i)}\right)^{L-1},\ \forall i\in[p],\ \bm{A}_{p+1}:=\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(p+1)}\bm{\Sigma}_{l+1}^{(p+1)^{T}},$ (54)
$\textstyle\bm{B}_{i}:=\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1},\ \forall i\in[p],\ \bm{B}_{p+1}:=\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p+1)^{T}}\bm{T}_{l+1}^{(p+1)},$ (55)
$c_{2}:=\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3y_{1}L}{2\sqrt{\lambda}\delta_{\sigma}\sigma^{*}_{\min}}+c_{1}+y_{1}p\sqrt{\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}$
$\qquad\ \left(\frac{L^{2}\eta_{1}}{2\sigma^{*}_{\min}}+\frac{3\sqrt{2g_{\max}}L^{2}\sigma_{\max}^{*}}{2\lambda\sigma_{\min}^{*}}\right).$ (57)
Then, we have
$\left\|(\bm{\Sigma}_{1}\bm{\Sigma}_{1}^{T})^{L-1}\bm{\Sigma}_{1}+\lambda\bm{\Sigma}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{A}_{p+1}\right)\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\right\|_{F}\leq c_{2}\|\nabla G(\bm{W})\|_{F},$ (58)
$\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}\leq c_{2}\|\nabla G(\bm{W})\|_{F}.$ (59)
Proof.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. It follows from Lemma 2 and 2 that (35)-(37) and (41) holds. Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ for each $l\in[L]$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$ for each $l=2,\dots,L$. This, together with (48), implies
@@
where $c_{1}$ is defined in (42). Using the SVD in (3.2.1), we have
$\textstyle\left\|(\bm{W}_{L}\bm{W}^{T}_{L})^{L-1}\bm{W}_{L}-\sqrt{\lambda}\bm{Y}\bm{W}_{L-1:1}^{T}+\lambda\bm{W}_{L}\right\|_{F}=\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}+\lambda\bm{\Sigma}_{L}\right\|_{F}$
$\geq$ $\textstyle\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}(\bm{H}_{2}-\bm{T}_{2})\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}$
$\textstyle-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}\bm{T}_{2}\bm{\Sigma}_{2}^{T}(\bm{H}_{3}-\bm{T}_{3})\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}-\dots-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\left(\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right)\bm{\Sigma}_{L-1}^{T}(\bm{H}_{L}-\bm{T}_{L})\right\|_{F}$
@@
where the first inequality uses the triangle inequality and the last inequality follows from (35), $\|\bm{H}_{l}\|=\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (60). This, together with (61), implies
$\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}$ (62)
@@
Using the diagonal structure of $\bm{\Sigma}_{l}$ in (3.2.1) and $\bm{T}_{l}=\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$, we compute
$\textstyle\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}=\mathrm{BlkD}\left(\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(1)}\bm{T}_{l+1}^{(1)},\dots,\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p)}\bm{T}_{l+1}^{(p)},\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p+1)}\bm{T}_{l+1}^{(p+1)}\right).$
@@
$\left\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\right\|_{F}^{2}$ $\textstyle=\sum_{j=1}^{g_{i}}\left(\sigma_{j}(\bm{\Sigma}_{l}^{(i)})-\sigma_{j}(\bm{\Sigma}_{L}^{(i)})\right)^{2}=\sum_{j=1}^{g_{i}}\left(\sum_{k=l}^{L-1}\left(\sigma_{j}(\bm{\Sigma}_{k}^{(i)})-\sigma_{j}(\bm{\Sigma}_{k+1}^{(i)})\right)\right)^{2}$
$\leq g_{i}(L-1)^{2}\left(\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}\right)^{2}.$
This further implies
$\left\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\right\|_{F}\leq\frac{3\sqrt{2g_{\max}}L\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}.$ (64)
Now we claim that it holds for each $l\in[L]$ and $i\in[p]$ that
$\left\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\right\|_{F}\leq\frac{\eta_{1}}{\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},$ (65)
where $\eta_{1}$ is defined in (56). To maintain the flow of the main proof, the proof of this claim is deferred to the end. Using the triangle inequality, we obtain
$\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\ \sum_{j=2}^{L}\|\underbrace{\textstyle\left(\prod_{l=1}^{L-j}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\!\!-\!\!\left(\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\right)\bm{\Sigma}_{L}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{j-2}}_{\bm{R}_{j}}\!\!\!\!\|_{F}.$ (66)
@@
[…]
where the first inequality uses (i) of Lemma 2 and $\|\bm{T}_{l+1}^{(i)}\|=1$ for all $l$ and $i$, and the second one is due the triangle inequality. This, together with (3.2.1), implies
@@
$\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)(\bm{\Sigma}_{L}^{(i)})^{L-1}\right\|_{F}+\left\|(\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{2}^{(i)}\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+$
$\textstyle\ \left\|\bm{\Sigma}_{L}^{(i)}\bm{T}_{2}^{(i)}(\bm{\Sigma}_{2}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{3}^{(i)}\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+\cdots+\left\|\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}(\bm{\Sigma}_{L-1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{L}^{(i)}\right\|_{F}$
@@
where the first inequality uses the triangle inequality, and the last inequality follows from (67), (35) in Lemma 2, $\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (64). This, together with (55), yields
$\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
@@
$\ \left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
$\leq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}$
$\quad+\sqrt{\lambda}y_{1}\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
@@
where $c_{2}$ is defined in (57). In (61), setting $l=1$ and applying the same argument, we obtain (58).
The rest of the proof is devoted to proving the claim (65). According to (3.2.1) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l=2,\dots,L$, we have for each $i\in[p]$,
$\left\|\bm{W}_{l}^{T}\bm{W}_{l}-\bm{W}_{l-1}\bm{W}_{l-1}^{T}\right\|_{F}=\left\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{H}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{T}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\right\|^{2}\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{l}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\right\|^{2}\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}-\left\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l}^{(i)})^{2}\right\|_{F}-\left\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\right\|_{F}$
[…]
where the second inequality follows from $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$. This together with (36), (60), and (64), yields that for each $l=2,\dots,L$ and $i\in[p]$, we have
$\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}\leq\eta_{1}\left\|\nabla G(\bm{W})\right\|_{F}.$ (69)